OpenStudy (elenathehomeschooler):

Complete the following for each system that involves one or more quadratic equations: (a) Solve the system algebraically 7y^2-5x^2+20x=3 5x^2+21y^2=209

OpenStudy (anonymous):

5x^2+21y^2-3(7y^2-5x^2+20x)=209-3*3 5x^2+21y^2-21y^2+15x^2-60x=200 20x^2-60x=200 x^2-3x-10=0 x^2-5x+2x-10=0 find x and then corresponding y